# The $r$-matrix structure on the moduli space of framed Higgs pairs

### v1

@@ 1 Introduction
In the elliptic (and trigonometric) case the $r$-matrix structure (1.1) was defined in [23, 8]; for the three cases, rational, trigonometric and elliptic, in fact there are families of pencils of compatible structures, as shown in [12]. The question, however, of a proper $r$-matrix structure for Hitchin’s systems over a curve of arbitrary genus $g$ remains open to this day and it is precisely the goal of this paper. In fact in [14] the authors conclude “[…] we can see that the Poisson geometry of rational surfaces suggests quite strongly that there is no nice Poisson extension of the Sklyanin bracket to arbitrary base curves”.
@@
$\left\{\displaystyle\mathop{\Phi}^{1}(w)\displaystyle\mathop{,}^{\otimes}\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\right\}\$$=\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,w)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,q)\displaystyle\mathop{\mathbf{K}}^{1}(q,r)\Pi,$
@@ 3 The higher genus $r$-matrix structure
Note the order of arguments of $\mathbf{K}$ which is dictated by the fact that the left and right sides should be bi-differentials. It is tempting to use the exact same expression but with the new understanding of $\Phi$ as a framed Higgs field and $\mathbf{K}$ as the non-abelian Cauchy kernel:
$\big{\{}\Phi_{ab}(p),\Phi_{cd}(q)\big{\}}=\Phi_{ad}(p)\mathbf{K}_{cb}(q,p)-\Phi_{ad}(q)\mathbf{K}_{cb}(p,q)-\Phi_{cb}(p)\mathbf{K}_{ad}(q,p)+\Phi_{cb}(q)\mathbf{K}_{ad}(p,q).$ (3.3)
However this is certainly not correct as we explain presently. Let us see what structure the pole at $p=t_{j}$ should have; this must start with a double pole with leading singular part $({\mathbf{v}}_{j}{\mathbf{h}}_{j}^{t})_{ab}$ because $t_{j}$ is coupled with the eigenvalue $\varkappa_{j}$. The first term has this structure in the $(a,b)$ indices, but there is also the third term with a double pole and the leading coefficient is $({\mathbf{f}}_{j}(q){\mathbf{h}}_{j}^{t})_{ad}$ which is incorrect. Instead we will prove
@@
$\Big{\{}\Phi_{ab}(w),\mathbf{K}_{cd}(q,r)\Big{\}}={\mathbf{K}_{ad}(w,r)\mathbf{K}_{cb}(q,r)-\mathbf{K}_{ad}(w,r)\mathbf{K}_{cb}(q,w)-\delta_{cb}\big{(}\mathbf{K}(w,q)\mathbf{K}(q,r)\big{)}_{ad}},$ (3.5)
@@
$\left\{\displaystyle\mathop{\Phi}^{1}(w)\displaystyle\mathop{,}^{\otimes}\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\right\}\$$=\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,w)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,q)\displaystyle\mathop{\mathbf{K}}^{1}(q,r)\Pi,$ (3.7)
@@
It is however rather remarkable, in that a direct verification appears significantly complicated. The proof in Section A is not particularly elegant but it is straightforward, relying heavily on the properties of the non-abelian Cauchy kernel.
@@ A.1 Poisson brackets of the kernel with $\Phi$
$=\sum_{j}\mathop{\mathrm{res}}\limits_{x=t_{j}}\left({\mathbf{K}_{c\ell}(q,x)\mathbf{K}_{ad}(p,x)-\mathbf{K}_{c\ell}(q,x)\mathbf{K}_{ad}(p,q)-\delta_{ad}\big{(}\mathbf{K}(q,p)\mathbf{K}(p,x)\big{)}_{c\ell}}\right)\left[\frac{N^{(j)}_{\ell b}}{x_{j}-\zeta_{j}}+\varkappa_{j}\delta_{\ell b}\right]\mathrm{d}x_{j}+$ (A.41)
@@ Appendix B Alternative construction of the Cauchy kernel
The formula (2.28) seems to require the computation of the period matrix, which is a transcendendal computation. This is unnecessary and one can choose differentials and a third kind differential that require only algebraic manipulations. Fix a non-special divisor $\mathscr{D}=\sum_{j=1}^{g}d_{j}$. There is a unique (scalar) Cauchy kernel $C(p,q)$ with divisor properties
@@
$C(p,q)=\left(\frac{E(z,y)}{(x-z)(y-w)}-\frac{E(x_{\infty},y)}{(x-x_{\infty})(y-y_{\infty})}+Q((x,y);(z,w))\right)\frac{\mathrm{d}x}{E_{x}(x,y)}$ (B.3)
where $Q(x,z;z,w)$ is a polyomial in $x,y$ with powers in the Newton polygon $P$ (thus corresponding to a holomorphic differential of $p$ and containing $g$ terms of the form $x^{n_{j}}y^{m_{j}}$ with $(n_{j},m_{j})\in P\subset{\mathbb{N}}\times{\mathbb{N}}$, $j=1,\dots g$) obtained by imposing the interpolating condition that the whole bracket vanishes at $(x,y)=(x_{j},y_{j})$, $j=1,\dots g$. Note also that this $C(p,q)$ is, in fact, the $n=1$ version of the non-abelian Cauchy kernel, and $\mathscr{D}$ is a the Tyurin divisor for a line bundle.

### v3

@@ 1 Introduction
In the elliptic (and trigonometric) case the $r$-matrix structure (1.1) was defined in [24, 9]; for the three cases, rational, trigonometric and elliptic, in fact there are families of pencils of compatible structures, as shown in [13]. The question, however, of a proper $r$-matrix structure for Hitchin’s systems over a curve of arbitrary genus $g$ is precisely the goal of this paper. In fact in [15] the authors conclude “[…] we can see that the Poisson geometry of rational surfaces suggests quite strongly that there is no nice Poisson extension of the Sklyanin bracket to arbitrary base curves”. The use of the Tyurin parametrization and a partial description of the rr–matrix was considered in [6] and hence the present paper does not appear in a vacuum. We comment below in Remark 1.2 on the added value of the present paper.
@@
$\left\{\displaystyle\mathop{\Phi}^{1}(w)\displaystyle\mathop{,}^{\otimes}\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\right\}\$$=\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,w)\Pi+\displaystyle\mathop{\mathbf{K}}^{1}(w,q)\displaystyle\mathop{\mathbf{K}}^{1}(q,r)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\Pi,$
@@
Example 1.1
Consider the expression ${\cal H}_{1}(q):=\mathrm{Tr}\Phi(q)$; this is a holomorphic differential (all residue at $t_{j}$, $\infty$ vanish) and a simple computation from (1.3) yields:
$\{\Phi(p),{\cal H}_{1}(q)\}=\left[\Phi(p),\mathbf{K}(q,p)\right].$ (1.4)
Note that in genus 0 this is automatically zero since then $\mathbf{K}$ is multiple of the identity matrix. The periods of $\mathrm{Tr}\Phi$ along a Lagrangian basis in homology forms a collection of $g$ commuting and independent hamiltonian functions and so the corresponding flows are given by the commutator of $\Phi$ with the corresponding matrix of periods of the Cauchy kernel, $M_{\gamma}(q)=\oint_{p\in\gamma}\mathbf{K}(p,q)$.
Similarly, consider ${\cal H}_{2}(q):=\frac{1}{2}\mathrm{Tr}\Phi^{2}(q)$. This a quadratic differential on $\mathcal{C}$ with a single double pole at $\infty$, in general, for a framed Higgs field (there are no poles at the Tyurin points). It generates certain Hamiltonians for the Hitchin system. Again, a simple computation using (1.3) yields
$\{\Phi(p),{\cal H}(q)\}=\left[\Phi(p),\Phi(q)\mathbf{K}(q,p)\right].$ (1.5)
Remark 1.2
After completing this work the author became aware of the very closely related [6]. In fact, modulo notational differences, the setup of [6] is precisely the same as the present one. The author there presents an $r$-matrix which has the same form as the one we give in Theorem 3.1 with two relatively important practical differences:
  1. 1.
    the rr–matrix’s existence is stated in Lemma 1, Lemma 3 ibidem: these are existence statements of what we call here the non-abelian Cauchy kernel. However the reference did not provide an explicit formula comparable to (2.28) below or the other explicit formulas contained in [3] for arbitrary Tyurin data.
  2. 2.
    while it was stated that the rr–matrix is dynamical, the Poisson brackets with the Higgs field were not evaluated, which is necessary to close the Poisson algebra. This was in fact recognized in Section 4 ibidem as a missing feature.
The present paper provides a fully explicit expression for the rr–matrix and also a complete description of the resulting Poisson algebra. See also App. B $\blacktriangle$
@@ 3 The higher genus $r$-matrix structure
Note the order of arguments of $\mathbf{K}$ which is dictated by the fact that the left and right sides should be bi-differentials. It is tempting to use the exact same expression (3.2) but with the new understanding of $\Phi$ as a framed Higgs field and $\mathbf{K}$ as the non-abelian Cauchy kernel. However this is certainly not correct as we explain presently. Let us see what structure the pole at $p=t_{j}$ should have; this must start with a double pole with leading singular part $({\mathbf{v}}_{j}{\mathbf{h}}_{j}^{t})_{ab}$ because $t_{j}$ is coupled with the eigenvalue $\varkappa_{j}$. The first term has this structure in the $(a,b)$ indices, but there is also the third term with a double pole and the leading coefficient is $({\mathbf{f}}_{j}(q){\mathbf{h}}_{j}^{t})_{ad}$ which is incorrect. Instead we will prove
@@
$\Big\{\Phi_{ab}(w),\mathbf{K}_{cd}(q,r)\Big\}={\mathbf{K}_{ad}(w,r)\mathbf{K}_{cb}(q,w)+\delta_{cb}\big(\mathbf{K}(w,q)\mathbf{K}(q,r)\big)_{ad}-\mathbf{K}_{ad}(w,r)\mathbf{K}_{cb}(q,r)},$ (3.4)
@@
$\left\{\displaystyle\mathop{\Phi}^{1}(w)\displaystyle\mathop{,}^{\otimes}\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\right\}\$$=+\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,w)\Pi+\displaystyle\mathop{\mathbf{K}}^{1}(w,q)\displaystyle\mathop{\mathbf{K}}^{1}(q,r)\Pi-\displaystyle\mathop{\mathbf{K}}^{1}(w,r)\displaystyle\mathop{\mathbf{K}}^{2}(q,r)\Pi,$ (3.6)
@@
It is however rather remarkable, in that a direct verification appears significantly complicated. In view of this difficulty it is not surprising then that the computation of the Poisson brackets $\{\Phi,\mathbf{K}\}$ was not completed in [6]. The proof in Section A is not particularly elegant but it is straightforward, relying heavily on the properties of the non-abelian Cauchy kernel.
@@ A.1 Poisson brackets of the kernel with $\Phi$
$=\sum_{j}\mathop{\mathrm{res}}\limits_{x=t_{j}}\left({\mathbf{K}_{c\ell}(q,x)\mathbf{K}_{ad}(p,q)+\delta_{ad}\big(\mathbf{K}(q,p)\mathbf{K}(p,x)\big)_{c\ell}-\mathbf{K}_{c\ell}(q,x)\mathbf{K}_{ad}(p,x)}\right)\left[\frac{N^{(j)}_{\ell b}}{x_{j}-\zeta_{j}}+\varkappa_{j}\delta_{\ell b}\right]\mathrm{d}x_{j}+$ (A.41)
@@ Appendix B Alternative construction of the Cauchy kernel
The formula (2.28) seems to require the computation of the period matrix, which is a transcendental computation. This is unnecessary and one can choose differentials and a third kind differential that require only algebraic manipulations. Fix a non-special divisor $\mathscr{D}=\sum_{j=1}^{g}d_{j}$. There is a unique (scalar) Cauchy kernel $C(p,q)$ with divisor properties
@@
$C(p,q)=\left(\frac{E(z,y)}{(x-z)(y-w)}-\frac{E(x_{\infty},y)}{(x-x_{\infty})(y-y_{\infty})}+Q((x,y);(z,w))\right)\frac{\mathrm{d}x}{E_{y}(x,y)}$ (B.3)
where $Q((x,z);(z,w))$ is a polyomial in $x,y$ with powers in the Newton polygon $P$ (thus corresponding to a holomorphic differential of $p$ and containing $g$ terms of the form $x^{n_{j}}y^{m_{j}}$ with $(n_{j},m_{j})\in P\subset{\mathbb{N}}\times{\mathbb{N}}$, $j=1,\dots g$) obtained by imposing the interpolating condition that the whole bracket vanishes at $(x,y)=(x_{j},y_{j})$, $j=1,\dots g$. Note also that this $C(p,q)$ is, in fact, the $n=1$ version of the non-abelian Cauchy kernel, and $\mathscr{D}$ is a the Tyurin divisor for a line bundle.